# One-loop QCD corrections to SSA in unweighted Drell-Yan processes

### v1

@@ VI.3 SFP contributions
In this work, we concentrate on the contribution of $\bar{q}\otimes T_{F}$. For this case, the diagrams giving SFP are shown in Fig.7. Mirror diagrams are not shown, which do not contributeEguchi et al. (2007).
@@
$\{I\langle P_{7}\rangle,I\langle L\rangle\}=$ $2\times\frac{\alpha_{s}\{1,l_{\perp,cs}\cdot\tilde{s}_{\perp}\cos\theta\}}{128\pi N_{c}^{2}}\bar{A}_{\epsilon}\int_{r}\bar{q}(x_{b})T_{F}(0,x_{2})\frac{g_{3}}{(1-\hat{x}_{2})_{+}(1-\hat{x}_{b})_{+}}.$ (132)
@@
So far, we have listed all unsubtracted hard coefficients relevant to $\bar{q}\otimes T_{F}$. For $I\langle P_{7}\rangle$, all results are the same as our previous resultChen et al. (2017).
@@ VII Renormalization and Subtraction
It is for $g_{s}^{B}\bar{M}_{B}^{(1)}$ that is renormalized as a whole,
@@
Next, we should make a collinear subtraction to remove the collinear divergence in the hard partCollins and Rogers (2008). The subtraction term is obtained by following replacement in tree level cross section,
@@
$\Delta\bar{q}(x)=$ $-\frac{\alpha_{s}}{2\pi}\Big{(}\frac{2}{\epsilon}-\gamma_{E}+\ln 4\pi\Big{)}\int_{x}^{1}\frac{dz}{z}P_{qq}(z)\bar{q}(\frac{x}{z}),$
@@
The kernel for $T_{F}$ is a little complicatedKang and Qiu (2009); Braun et al. (2009); Zhou et al. (2010); Ma and Wang (2012); Schafer and Zhou (2012); Kang and Qiu (2012), which reads
@@ VII.1 $I\langle P_{7}\rangle$ subtraction
Complete real corrections can be read from eqs.(121,127,130,133). Here we list their divergent parts for convenience.
@@ VII.2 $I\langle L\rangle$ subtraction
However, virtual corrections do not have the same relation. We have
$I\langle L\rangle\Big{|}_{v}=$ $l_{\perp,cs}\cdot\tilde{s}_{\perp}\cos\theta 2\times\frac{\alpha_{s}}{128\pi N_{c}^{2}}\bar{A}_{\epsilon}\int_{\tau}^{1}\frac{d\xi}{\xi}\int_{\tau_{\xi}}^{1}\frac{dx_{b}}{x_{b}}\int\frac{dx_{2}}{x_{2}}\Big{\{}\bar{q}(x_{b})T_{F}(x_{2},x_{2})\Big{[}\delta(1-\hat{x}_{2})\delta(1-\hat{x}_{b})g_{0}\Big{]}$
$+\bar{q}(x_{b})T_{F}(x_{2}^{*},x_{2}^{*}-x_{2})\delta(1-\hat{x}_{b})\Big{[}g_{1}^{(c)}\theta(x_{2}-x_{2}^{*})+\text{finite.}\Big{]}\Big{\}},$ (166)
@@
$g_{0}=$ $-32(N_{c}^{2}-1)\Big{(}\frac{8}{\epsilon^{2}}+\frac{6}{\epsilon}+8-\pi^{2}\Big{)},$
$g_{1}^{(c)}=$ $128\hat{x}_{2}(\frac{1}{\epsilon}+1).$ (167)
$g_{1}^{(c)}$ is given by nonpole contributions. Other finite terms can be obtained from eq.(99). The boundary for $x_{2}$ must be made clear here. Because of different theta functions, the divergent part and finite part have different boundaries. For divergent part, the theta function is $\theta(x_{2}-x^{*}_{2})$, so, $x_{2}>x^{*}_{2}=\xi>0$. On the other hand, for $T_{F}(x_{1},x_{2})$, the support is $|x_{1}|<1$, $|x_{2}|<1$ and $|x_{1}-x_{2}|<1$, so, for $T_{F}(x^{*}_{2},x^{*}_{2}-x_{2})$, we have $|x_{2}|<1$. Together with $x_{2}>\xi$, we have $\xi<x_{2}<1$. Keeping the divergent part only, we have
$I\langle L\rangle\Big{|}_{v}\doteq$ $l_{\perp,cs}\cdot\tilde{s}_{\perp}\cos\theta 2\times\frac{\alpha_{s}}{128\pi N_{c}^{2}}\bar{A}_{\epsilon}\int_{r}\Big{\{}\bar{q}(x_{b})T_{F}(x_{2},x_{2})\Big{[}\delta(1-\hat{x}_{2})\delta(1-\hat{x}_{b})g_{0}\Big{]}$
@@
where $\int_{r}$ is defined by eq.(120). Without $g_{1}^{(c)}$, above result is 4-4 times of $I\langle P_{7}\rangle|_{v}$.
The real corrections of $I\langle L\rangle$ are given by eqs.(123, 128,131,134). The divergent part is summarized as follows:
$I\langle L\rangle\Big{|}_{r}\doteq$ $2\times\frac{\alpha_{s}}{128\pi N_{c}^{2}}\bar{A}_{\epsilon}\int_{r}\Big{\{}\bar{q}(x_{b})T_{F}(x_{2},x_{2})\Big{[}\delta(1-\hat{x}_{2})\delta(1-\hat{x}_{b})g_{0}^{(a)}+\delta(1-\hat{x}_{b})g_{1}^{(a)}+\delta(1-\hat{x}_{2})g_{2}^{(a)}\Big{]}$
@@
$g_{0}|_{r+v}=$ $g_{0}|_{v}+g_{0}|_{r}^{a+b+c}=-\frac{192(N_{c}^{2}-1)}{\epsilon}+32(N_{c}^{2}-1)(\pi^{2}-8).$ (173)
However, the subtraction term is
@@
So,
$g_{0}|_{r+v}-g_{0}|_{sub}=-\frac{128N_{c}^{2}}{\epsilon}.$ (175)
This is the divergence that cannot be subtracted!
Another problem is for the non-pole virtual correction. In the above we have written this part into $g_{1}^{(c)}$, i.e.,
$I\langle L\rangle\Big{|}_{v}\supset$ $\text{const.}\int_{r}\bar{q}(x_{b})T_{F}(x_{2}^{*},x_{2}^{*}-x_{2})\delta(1-\hat{x}_{b})\theta(x_{2}-x_{2}^{*})g_{1}^{(c)},\ g_{1}^{(c)}=128\hat{x}_{2}(\frac{1}{\epsilon}+1),$
$\text{const.}=$ $(l_{\perp,cs}\cdot\tilde{s}_{\perp}\cos\theta)2\times\frac{\alpha_{s}}{128\pi N_{c}^{2}}\bar{A}_{\epsilon}.$ (176)
We notice that when $\hat{x}_{b}=1$, $x_{2}^{*}=x_{1}^{*}$ with $x_{1}^{*}$ determined by the hard pole condition $(k_{1}-q)^{2}=0$. When $\hat{x}_{b}=1$, the magnitude of the transverse momentum of virtual photon is
$q_{t}=\sqrt{\frac{Q^{2}}{\hat{x}_{b}}(1-\hat{x}_{2})(1-\hat{x}_{b})}=0.$ (177)
Then,
$q^{-}=\frac{Q^{2}+q_{t}^{2}}{2q^{+}}=\frac{Q^{2}}{2\xi p_{a}^{+}}.$ (178)
So,
$x_{1}^{*}=\frac{Q^{2}}{2p_{a}\cdot q}=\xi.$ (179)
This is just $x_{2}^{*}$. Thus, non-pole contribution can be written as
$I\langle L\rangle\Big{|}_{v}\supset\text{const.}\int_{r}\bar{q}(x_{b})T_{F}(x_{1}^{*},x_{1}^{*}-x_{2})\delta(1-\hat{x}_{b})g_{1}^{(c)},$ (180)
and is actually a part of hard pole contribution. After adding it to corresponding real contribution, we find
$I\langle L\rangle\Big{|}_{v}+I\langle L\rangle\Big{|}_{r}\supset$ $\text{const.}\int_{r}\bar{q}(x_{b})T_{F}(x_{1}^{*},x_{1}^{*}-x_{2})\delta(1-\hat{x}_{b})\Big{[}-4\frac{16-32\hat{x}_{2}}{\epsilon}+128\hat{x}_{2}\Big{]}.$ (181)
@@
The divergence is the same as that of subtraction term,
$I\langle L\rangle\Big{|}_{sub}\supset-4\times\text{const.}\int_{r}\bar{q}(x_{b})T_{F}(x_{2}^{*},x_{2}^{*}-x_{2})\delta(1-\hat{x}_{b})\frac{16-32\hat{x}_{2}}{\epsilon}.$ (182)
Thus, the divergent non-pole part in virtual correction is safe, and should be viewed as a part of hard pole contribution.
Other $g_{i}$’s are also safe:
  1. 1.
    $g_{1}^{a}$:
    $g_{1}^{(a)}|_{r+v}=g_{1}|_{r}^{(a)}=\frac{64(1+\hat{x}_{2}^{2})}{\epsilon(1-\hat{x}_{2})_{+}},\ g_{1}^{(a)}|_{sub}=-4\times\frac{-16(1+\hat{x}_{2}^{2})}{\epsilon(1-\hat{x}_{2})_{+}},$ (183)
    so,
    $g_{1}^{(a)}|_{r+v}-g_{1}^{(a)}|_{sub}=0.$ (184)
  2. 2.
    $g_{1}^{(b)}$:
    $g_{1}|_{r+v}^{(b)}=g_{1}|_{r}^{(b)}=-\frac{64N_{c}^{2}(1+\hat{x}_{2})}{\epsilon(1-\hat{x}_{2})_{+}},\ g_{1}|_{sub}^{(b)}=-4\times\frac{16N_{c}^{2}(1+\hat{x}_{2})}{\epsilon(1-\hat{x}_{2})_{+}}.$ (185)
    So,
    $g_{1}|_{r+v}^{(b)}-\ g_{1}|_{sub}^{(b)}=0.$ (186)
  3. 3.
    $g_{2}$:
    $g_{2}|_{r+v}=g_{2}|_{r}^{(a+b+c)}=-\frac{64(N_{c}^{2}-1)(1+\hat{x}_{b}^{2})}{\epsilon(1-\hat{x}_{b})_{+}},\ g_{2}|_{sub}=-4\times\frac{16(N_{c}^{2}-1)(1+\hat{x}_{b}^{2})}{\epsilon(1-\hat{x}_{b})_{+}}.$ (187)
    So,
    $g_{2}|_{r+v}-g_{2}|_{sub}=0.$ (188)
In summary, all divergences can be removed by collinear subtraction, except for $g_{0}$-term, i.e.,
$I\langle L\rangle\Big{|}_{r+v}-I\langle L\rangle\Big{|}_{sub}=2\times\frac{\alpha_{s}l_{\perp,cs}\cdot\tilde{s}_{\perp}\cos\theta}{128\pi N_{c}^{2}}\bar{A}_{\epsilon}\int_{r}\bar{q}(x_{b})T_{F}(x_{2},x_{2})\Big{[}\delta(1-\hat{x}_{2})\delta(1-\hat{x}_{b})\frac{-128N_{c}^{2}}{\epsilon}\Big{]}+\text{finite.}$ (189)
This indicates twist-3 collinear factorization does not hold. In $\cdots$], $N_{c}^{2}$ should be understood as $C_{A}N_{c}$, and this term comes from Fig.5(c). As an example, we present the calculation of HP contribution of Fig.5(c) in App.B. More clearly, we redraw the diagrams in Fig.8. We analyze the divergence of these two diagrams in next section with eikonal approximation.
VIII Eikonal approximation for HP contribution
From the result of last section, it is the divergence from HP breaks the factorization. The divergence is proportional to $\delta(1-\hat{x}_{2})\delta(1-\hat{x}_{b})$, which indicates the final gluon with momentum $k_{g}$ is a soft gluon. The two diagrams having HP contribution proportional to $C_{A}N_{c}$ are shown in Fig.8.
Figure 8: The diagrams having HP contribution proportional to $C_{A}N_{c}$.
The contribution proportional to $\delta(1-\hat{x}_{2})\delta(1-\hat{x}_{b})$ can be given by soft approximation or eikonal approximation. The final gluon in this approximation is a soft one, with $k_{g}^{\pm}\sim k_{g\perp}\sim\lambda$. $\lambda$ is a small quantity. The hard pole condition is $(k_{1}-q)^{2}=0$ or $(k_{b}+k-k_{g})^{2}=0$. So,
$2k_{b}^{-}(k^{+}-k_{g}^{+})+k_{g\perp}^{2}=0.$ (190)
Because $k_{g\perp}\sim\lambda$, we must have
$k^{+}-k_{g}^{+}\sim\lambda^{2},$ (191)
which also implies $k^{+}\sim\lambda$. This is different from twist-2 case, where all external momenta are $O(1)$.
To simplify the analysis, we use $G_{\perp}$ as the initial gluon from polarized hadron to do collinear expansion. Then,
$I\langle L\rangle=$ $\int d^{n}q\delta(q^{2}-Q^{2})L_{\mu\nu}W^{\mu\nu}$
$=$ $\int\frac{dq^{+}}{2q^{+}}\int d^{n-2}q_{\perp}\int dk_{2}^{+}dk^{+}dk_{b}^{-}\bar{q}(x_{b})M^{(1)}_{\gamma^{+},\partial^{+}G_{\perp}}\tilde{s}_{\perp}^{\rho}\frac{i}{k^{+}}(-i\pi)\delta((k_{1}-q)^{2})\hat{H}_{\rho}^{\mu\nu}L_{\mu\nu},$ (192)
where
$\hat{H}_{\rho}^{\mu\nu}=$ $\int\frac{d^{n}k_{g}}{(2\pi)^{n}}2\pi\delta_{+}(k_{g}^{2})\delta^{n}(k_{2}+k_{b}-q-k_{g})H_{1,2}^{\mu\nu,\rho}$
$=$ $\frac{1}{(2\pi)^{n-1}}\hat{x}_{2}\delta(q_{t}^{2}-\frac{Q^{2}}{\hat{x}_{2}}(1-\hat{x}_{2})(1-\hat{x}_{b}))H_{1,2}^{\mu\nu,\rho}.$ (193)
$H_{1,2}^{\mu\nu,\rho}$ is the hard part for Fig.8(a) and (b), respectively. The explicit expressions are
[…]
with $\Gamma_{\rho\tau\lambda}(k,k_{g}-k,-k_{g})=g_{\rho\tau}(2k-k_{g})_{\lambda}+g_{\tau\lambda}(2k_{g}-k)_{\rho}+g_{\lambda\rho}(-k_{g}-k)_{\tau}$. After integrating out $q_{\perp}$ and $k^{+}$, we get
$I\langle L\rangle=$ $\frac{\pi\Omega_{n-2}}{4(2\pi)^{n-1}}\int\frac{dq^{+}}{q^{+}}\int dk_{2}^{+}dk_{b}^{-}\bar{q}(x_{b})M^{(1)}_{\gamma^{+},\partial^{+}G_{\perp}}\tilde{s}_{\perp\rho}(q_{t}^{2})^{-\epsilon/2}\frac{1}{2k^{+}q^{-}}H_{1,2}^{\mu\nu,\rho}L_{\mu\nu},\ q_{t}^{2}=\frac{Q^{2}}{\hat{x}_{2}}(1-\hat{x}_{2})(1-\hat{x}_{b}).$ (195)
The leading power behavior of the integrand is
$\frac{1}{2k^{+}q^{-}}H_{1,2}^{\mu\nu,\rho}L_{\mu\nu}\sim\frac{1}{q_{\perp}^{2}}\sim\frac{1}{\lambda^{2}}.$ (196)
After expansion in $\epsilon$, this part gives $\delta(1-\hat{x}_{2})\delta(1-\hat{x}_{b})$. Keeping leading power contribution only, we get
$I\langle L\rangle=$ $\text{K.}\int\frac{dq^{+}}{q^{+}}\int dk_{2}^{+}dk_{b}^{-}\bar{q}(x_{b})T_{F}(x_{1}^{*},x_{2})\tilde{s}_{\perp}^{\rho}(q_{t}^{2})^{-\epsilon/2}\Big{[}\frac{N_{1}^{\rho}}{2k\cdot q[(k_{b}-k_{g})^{2}][(k_{g}-k)^{2}]}+\frac{N_{2}^{\rho}\Big{|}_{\text{II}}+N_{2}^{\rho}\Big{|}_{\text{III}}}{2k\cdot q[(k-k_{g})^{2}][(k_{2}-k_{g})^{2}]}\Big{]},$
$\text{K.}=$ $\frac{\Omega_{n-2}}{4(2\pi)^{n-1}}\frac{[-ig_{s}^{2}f^{abc}Tr(T^{c}T^{b}T^{a})]}{8N_{c}^{2}C_{F}}.$ (197)
Because $^{+}\sim\lambda, k+k$$^{+}-k_{g}^{+}\sim\lambda^{2}$, the two denominators are of order $\lambda^{4}$. $N_{1}^{\rho}$ is given by Fig.8(a) and $N_{2}^{\rho}|_{\text{II}}$, $N_{2}^{\rho}|_{\text{III}}$ are given by Fig.8(b). Their expressions are
$N_{1}^{\rho}=$ $-16k_{g}\cdot k_{b}[k^{\nu}g_{\perp}^{\mu\rho}+k^{\mu}g_{\perp}^{\nu\rho}]L_{\mu\nu},$
$N_{2}^{\rho}|_{\text{II}}=$ $\frac{32q_{\perp}^{2}}{n-2}L_{\mu\nu}\Big{[}-\Big{(}\frac{\partial}{\partial q_{\perp}}\Big{)}_{L}g_{\perp}^{\mu\nu}+[k_{2}^{\mu}g_{\perp}^{\nu\rho}+k_{2}^{\nu}g_{\perp}^{\mu\rho}]+\frac{1}{2}\Big{(}[k_{b}^{\nu}g_{\perp}^{\mu\rho}+k_{b}^{\mu}g_{\perp}^{\nu\rho}]-[k_{2}^{\mu}g_{\perp}^{\nu\rho}+k_{2}^{\nu}g_{\perp}^{\mu\rho}]\Big{)}\Big{]},$
$N_{2}^{\rho}|_{\text{III}}=$ $16k\cdot k_{g}[k_{b}^{\mu}g_{\perp}^{\nu\rho}+k_{b}^{\nu}g_{\perp}^{\mu\rho}]L_{\mu\nu}.$ (198)
All of them are of order $\lambda^{2}$. In the above, $L^{\mu\nu}$ has been expanded in $q_{\perp}$ at $q_{\perp}=0$. The derivative $(\partial/\partial q_{\perp}^{\rho})_{L}$ in $N_{2}^{\rho}$ indicates the derivative acts on $L^{\mu\nu}$. After taking the derivative, $q_{\perp}$ is set to zero. The benefit of this decomposition is $N_{2}^{\rho}|_{\text{II}}$ alone satisfies QED gauge invariance, while $N_{1}^{\rho}$ and $N_{2}^{\rho}|_{\text{III}}$ together recovers such a symmetry. As can be seen,
$\frac{N_{1}^{\rho}}{2k\cdot q[(k_{b}-k_{g})^{2}][(k_{g}-k)^{2}]}+\frac{N_{2}^{\rho}\Big{|}_{\text{III}}}{2k\cdot q[(k-k_{g})^{2}][(k_{2}-k_{g})^{2}]}=\frac{4L_{\mu\nu}}{k_{2}\cdot qq_{\perp}^{2}}\Big{[}(k_{2}^{\mu}g_{\perp}^{\nu\rho}+k_{2}^{\nu}g_{\perp}^{\mu\rho})-(k_{b}^{\mu}g_{\perp}^{\nu\rho}+k_{b}^{\nu}g_{\perp}^{\mu\rho})\Big{]}.$ (199)
If we replace $L^{\mu\nu}$ by $q^{\mu}$ or $q^{\nu}$, the above is zero. This tensor also appears in $N_{2}|_{\text{II}}$. Then, the sum is
$I\langle L\rangle=$ $\text{K.}\int\frac{dq^{+}}{q^{+}}\int dk_{2}^{+}dk_{b}^{-}\bar{q}(x_{b})T_{F}(x_{1}^{*},x_{2})\tilde{s}_{\perp}^{\rho}(q_{t}^{2})^{-\epsilon/2}\frac{4L_{\mu\nu}}{k_{2}\cdot qq_{\perp}^{2}}\Big{\{}\frac{4}{2-\epsilon}\Big{[}-k_{2}\cdot k_{b}g_{\perp}^{\mu\nu}\Big{(}\frac{\partial}{\partial q_{\perp}^{\rho}}\Big{)}_{L}+(k_{2}^{\mu}g_{\perp}^{\nu\rho}+k_{2}^{\nu}g_{\perp}^{\mu\rho})\Big{]}$
$+(1-\frac{2}{2-\epsilon})\Big{[}(k_{2}^{\mu}g_{\perp}^{\nu\rho}+k_{2}^{\nu}g_{\perp}^{\mu\rho})-(k_{b}^{\mu}g_{\perp}^{\nu\rho}+k_{b}^{\nu}g_{\perp}^{\mu\rho})\Big{]}\Big{\}}.$ (200)
Because
$g_{\perp\mu\nu}\Big{(}\frac{\partial L^{\mu\nu}}{\partial q_{\perp}^{\rho}}\Big{)}_{q_{\perp}=0}=0,$
$L_{\mu\nu}(k_{2}^{\mu}g_{\perp}^{\nu\rho}+k_{2}^{\nu}g_{\perp}^{\mu\rho})\Big{|}_{q_{\perp}=0}=-L_{\mu\nu}(k_{b}^{\mu}g_{\perp}^{\nu\rho}+k_{b}^{\nu}g_{\perp}^{\mu\rho})\Big{|}_{q_{\perp}=0}=8k_{2}\cdot k_{b}\cos\theta l_{\perp,cs}^{\rho},$ (201)
we have

### v2

@@ VI.3 SFP contributions
In this work, we concentrate on the contribution of $\bar{q}\otimes T_{F}$. For this case, the diagrams giving SFP are shown in Fig.7. Mirror diagrams are not shown, which do not contribute[7].
@@
$\{I\langle P_{7}\rangle,I\langle L\rangle\}=$ $2\times\frac{\alpha_{s}\{1,l_{\perp,cs}\cdot\tilde{s}_{\perp}\cos\theta\}}{128\pi N_{c}^{2}}\bar{A}_{\epsilon}\int_{r}\bar{q}(x_{b})T_{F}(0,-x_{2})\frac{g_{3}}{(1-\hat{x}_{2})_{+}(1-\hat{x}_{b})_{+}}.$ (144)
@@
So far, we have listed all unsubtracted hard coefficients relevant to $\bar{q}\otimes T_{F}$. For $I\langle P_{7}\rangle$, all results are the same as our previous result[15].
@@ VII Renormalization and Subtraction
It is for $g_{s}^{B}\bar{M}_{B}^{(1)}$ that is renormalized as a whole quantity,
@@
Next, we should make a collinear subtraction to remove the collinear divergence in the hard part[36]. The subtraction term is obtained by following replacement in tree level cross section,
@@
$\Delta\bar{q}(x)=$ $-\frac{\alpha_{s}}{2\pi}\Big(\frac{2}{\epsilon}-\gamma_{E}+\ln 4\pi\Big)\int_{x}^{1}\frac{dz}{z}P_{qq}(z)\bar{q}(\frac{x}{z})\equiv-\frac{\alpha_{s}}{2\pi}\Big(\frac{2}{\epsilon}-\gamma_{E}+\ln 4\pi\Big)P_{qq}\otimes\bar{q}(x),$
@@
The kernel for $T_{F}$ is a little complicated[37, 38, 25, 39, 40, 41], which reads
@@ VII.1 $I\langle P_{7}\rangle$ subtraction
Complete real corrections can be read from eqs.(133,139,142,145). Here we list their divergent parts for convenience.
@@ VII.2 $I\langle L\rangle$ subtraction
However, virtual corrections do not have the same relation. The divergent part is
$I\langle L\rangle\Big|_{v}^{div}=$ $2\times l_{\perp,cs}\cdot\tilde{s}_{\perp}\cos\theta\frac{\alpha_{s}}{128\pi N_{c}^{2}}\bar{A}_{\epsilon}\int_{\tau}^{1}\frac{d\xi}{\xi}\int_{\tau_{\xi}}^{1}\frac{dx_{b}}{x_{b}}\int\frac{dx_{2}}{x_{2}}\Big\{\bar{q}(x_{b})T_{F}(x_{2},x_{2})\Big[\delta(1-\hat{x}_{2})\delta(1-\hat{x}_{b})g_{0}\Big]$
$+\frac{dx_{2}}{x_{2}}\bar{q}(x_{b})T_{F}(x_{2}^{*},x_{2}^{*}-x_{2})\delta(1-\hat{x}_{b})\Big[g_{1}^{(c)}\theta(x_{2}-x_{2}^{*})\Big]\Big\},$ (178)
@@
$g_{0}=$ $-32(N_{c}^{2}-1)\Big(\frac{8}{\epsilon^{2}}+\frac{6}{\epsilon}\Big)+\frac{128N_{c}^{2}}{\epsilon},$
$g_{1}^{(c)}=$ $\frac{128\hat{x}_{2}}{\epsilon}.$ (179)
The last term of $g_{0}$ comes from $\delta(x)$ part of $I\langle L\rangle|_{v}^{NP}$ in eq.(111), and another term comes from SGP contribution. $g_{1}^{(c)}$ is given by nonpole contributions. For $g_{1}^{(c)}$ term, the boundary of $x_{2}$ should be made explicit. Generally, the support of $T_{F}(x_{1},x_{2})$ is $|x_{1,2}|<1$, $|x_{1}-x_{2}|<1$. For $T_{F}(x_{2}^{*},x_{2}^{*}-x_{2})$, because $0<x_{2}^{*}<1$ the allowed region of $x_{2}$ is $x_{2}^{*}-1<x_{2}<1$. Considering $\theta(x_{2}-x_{2}^{*})$, we finally have $x_{2}^{*}<x_{2}<1$, which is precisely the same as that of hard pole contribution. Thus,
$I\langle L\rangle\Big|_{v}^{div}=$ $2\times l_{\perp,cs}\cdot\tilde{s}_{\perp}\cos\theta\frac{\alpha_{s}}{128\pi N_{c}^{2}}\bar{A}_{\epsilon}\int_{r}\Big\{\bar{q}(x_{b})T_{F}(x_{2},x_{2})\Big[\delta(1-\hat{x}_{2})\delta(1-\hat{x}_{b})g_{0}\Big]$
@@
where $\int_{r}$ is defined by eq.(132).
The real corrections of $I\langle L\rangle$ are given by eqs.(135, 140,143,146). The divergent part is summarized as follows:
$I\langle L\rangle\Big|_{r}^{div}=$ $2\times l_{\perp,cs}\cdot\tilde{s}_{\perp}\cos\theta\frac{\alpha_{s}}{128\pi N_{c}^{2}}\bar{A}_{\epsilon}\int_{r}\Big\{\bar{q}(x_{b})T_{F}(x_{2},x_{2})\Big[\delta(1-\hat{x}_{2})\delta(1-\hat{x}_{b})g_{0}^{(a)}+\delta(1-\hat{x}_{b})g_{1}^{(a)}+\delta(1-\hat{x}_{2})g_{2}^{(a)}\Big]$
@@
$g_{0}|_{r+v}=$ $g_{0}|_{v}+g_{0}|_{r}^{a+b+c}=-\frac{64N_{c}^{2}-192}{\epsilon}.$ (185)
This is the same as the corresponding subtraction term, which is
@@
Thus $g_{0}|_{r+v}-g_{0}|_{sub}$ is finite.
For $g_{1}^{(c)}$ parts in both virtual and real corrections, we notice that they are proportional to $\delta(\hat{x}_{b}-1)$. This implies
$q_{t}=\sqrt{\frac{Q^{2}}{\hat{x}_{b}}(1-\hat{x}_{2})(1-\hat{x}_{b})}=0,\ q^{-}=\frac{Q^{2}+q_{t}^{2}}{2q^{+}}=\frac{Q^{2}}{2\xi p_{a}^{+}},$ (187)
and
$x_{1}^{*}=\frac{Q^{2}}{2p_{a}^{+}q^{-}}=\xi=x_{2}^{*}.$ (188)
Thus, $g_{1}^{(c)}$ part in virtual correction can be understood as a hard pole contribution. This point is noticed also by [21]. Because of $T_{F}(x_{1},x_{2})=T_{F}(x_{2},x_{1})$, these two $g_{1}^{(c)}$ can be added together. The result is
$g_{1}^{(c)}|_{v}+g_{1}^{(c)}|_{r}=\frac{128\hat{x}_{2}-64}{\epsilon}.$ (189)
@@
This is the same as the divergence in corresponding subtraction term $g_{1}^{(c)}|_{sub}$. Thus, $g_{1}^{(c)}|_{r+v}-g_{1}^{(c)}|_{sub}$ is finite.
Other divergences in real correction, including $g_{1}^{(a)}|_{r}$, $g_{1}^{(b)}|_{r}$ and $g_{2}|_{r}$, are also checked. They are the same as the divergences of corresponding subtraction terms.
In summary, for $I\langle L\rangle$, all divergences in real and virtual corrections can be subtracted out by subtraction terms constructed from evolution equation of $T_{F}(x,x)$. Thus, twist-3 factorization holds at one-loop level for SSA of DY. A new feature appearing here is the divergent non-pole part of virtual correction has the same form as hard pole contribution of real corrections. This is highly nontrivial. The calculation performed in coordinate space at twist-3 level [42, 43] should be helpful to understand how the factorization appears. We will use this method to do calculation again in future.
VII.3 Final results
Because we have checked that all divergences from real and virtual corrections can be removed consistently by renormalization and subtraction, the final one-loop correction is finite and can be obtained from
$I\langle L\rangle|_{final}=I\langle L\rangle|_{r+v}-I\langle L\rangle|_{sub},$ (190)
by ignoring all divergences and setting $\bar{A}_{\epsilon}=1$. Results for $I\langle P_{7}\rangle$ are obtained in the same way and are the same as [15]. For simplicity, we just show the result of $I\langle L\rangle$ here. It is organized as